OpenStudy (anonymous):

once you rationalize the fraction: 1 / 11x sqrt 5 - 4y sqrt 3 what do you get?

OpenStudy (anonymous):

what is the fraction exactly ?

OpenStudy (anonymous):

im guessing u can use conjugate pairs?

OpenStudy (anonymous):

a headache :p

OpenStudy (anonymous):

\[1/ ( 11 x \sqrt{5} ) - 4 y \sqrt{3}\]

OpenStudy (anonymous):

is it 1/ the whole thing?

OpenStudy (anonymous):

Is it like this : \[\frac1{11x\sqrt5-4y\sqrt3}\] or : \[\frac1{11x\sqrt5}-4y\sqrt3\]??

OpenStudy (anonymous):

the second thing

OpenStudy (anonymous):

OK ! We have : \[\frac1{11x\sqrt5}-4y\sqrt3=\frac{1-44xy\sqrt{15}}{11x\sqrt5}=\frac{\sqrt5-220xy\sqrt3}{55x}\]